# Cross-Encoder Rediscovers a Semantic Variant of BM25

### v1

@@ 5.1. Formalizing the Hypothesized Function
  3. (3)
    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, analogous to the product of IDF and TF in BM25.
Since the formalization contains not only the hypothesized computation function but also $U_{0}$ and $MS$ from earlier subparts of the circuit, the linear model is an effective representation of the circuit.
5.2. Assessing the Linear Model Fit
@@
Given the linear representation, we can determine whether our hypothesis holds by comparing the linear model’s performance against the cross-encoder’s actual relevance scores.
@@
We first train a linear regression model using our base dataset. As the number of coefficients in our linear regression scales quickly with each additional included query term, we apply a query length cutoff to five tokens in this initial experiment to first establish feasibility of representing the circuit in this manner. When performing forward passes on the samples, we extract the following for each query token: (1) the MS (Matching Score), calculated as the sum of the query token’s attention over document tokens from the 13 Matching Heads, and (2) the token’s corresponding value in $U_{0}$, obtained by applying SVD to the embedding matrix. These serve as the input features (“x”) for the linear regression model, while the cross-encoder’s ranking scores are used as the target values (“y”). We then split the data into an 80/20 train-test set to evaluate how well the linear representation of the Semantic Scoring Circuit predicts the cross-encoder’s relevance scores.
@@
The linear regression model achieves a high pearson correaltion (corr = 0.8157, p ¡ 0.001) with ground-truth relevance scores, showing that it effectively captures the core components of the cross-encoder’s scoring mechanism in a simplified and interpretable form. This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p ¡ 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components more effectively capture the signals that the cross-encoder utilizes for ranking.
5.3. Generalization across IR datasets
We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept. Now, we ask the question, how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on? In this section, we extend our analysis to 12 different IR datasets [^1] Due to computational resources, we use a subset of BEIR datasets  (Thakur et al., 2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains. and investigate various query lengths.
Dataset. Since our linear model relies on fixed features based on query length, we stratify the dataset based on query length. To ensure sufficient samples, we aggregate samples from 12 datasets and only retain groups with more than 500 samples. We incorporate groups with query lengths from 1 to 22, which correponds to the 99.95th percentile of MSMARCO’s query length distribution to exclude outliers with excessively long queries that the model has rarely seen during training. This process results in a total of 19 groups, comprising 278194 samples, with an average group size of 14641.790 samples. While this approach may be unconventional for retrieval, our goal is to simply demonstrate consistency between the linear model’s relevance score and the cross-encoder to confirm the hypothesized function of the Relevance Scoring Heads.
Experiments and Results. We train 22 linear regression models, one for each query length (1-22) to predict the cross-encoder’s relevance scores. We evaluate each model using: (1) Pearson Correlation: quantifies the correlation between predicted and actual relevance scores; (2) Spearman Rank Correlation: assesses the consistency of the ranked lists between predictions and cross-encoder outputs; (3) NDCG@10: measures ranking performance and ensures no significant performance discrepancies. We include BM25 and a randomized set of linear regression features as baselines.
Table 2 shows the results, and we report median values to account for observable skewness caused by outliers in the data. We observe a Pearson correlation of ranking scores with a median of 0.8401 (median p ¡ 0.001), a Spearman rank correlation of 0.7619 (median p = 0.072), and an 88.4% alignment with cross-encoder performance in terms of NDCG@10.
The experimental results confirm the hypothesized function of the Relevance Scoring Heads (§4.1). Since this linear model summarizes the whole circuit as it is structured to incorporate both the computation and the necessary components, the high correlation with the cross-encoder’s performance shows that our circuit understanding has captured the core part of the cross-encoder’s relevance ranking mechanism.
Table 2. Comparison of Pearson, Spearman Rank, and NDCG@10 for SemanticBM and Baselines.
@@ 5.3. Generalization across IR datasets
6. Insights into Model Mechanisms
A two-stage process for query-document interactions. Our circuit analysis reveals that query-document interactions happen in two stages, bridging previous conflicting perspectives on how BERT cross-encoders form query representations. In the early to mid layers (0-8), matching heads attend to document terms that are exact or “soft” matches to query terms (§4.3). This first stage aligns with Qiao et al. (2019), who find that BERT focuses on exact query term matches or close paraphrases in the document, suggesting query-dependent document representations.
In mid to late layers (8-9), document representations are aggregated into query token representations by Contextual Query Representation heads (Section 4.2), consistent with Zhan et al. (2020) who argue that document tokens are largely query-independent. These two stages offer an explanation to previous conflicting perspectives on the role of query-document interactions: document representations initially have a query-specific relevance signal, but later, this signal is transferred to the query tokens. This dual-stage interaction process provides a more nuanced understanding of how query-document interactions evolve across cross-encoder layers.
Is (semantic) BM25 the whole story? While the Semantic Scoring circuit bears a strong resemblance to BM25, there also may be other factors influencing relevance computation that are out of the scope of investigation for this work. In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022). However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
Additionally, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores. There are a couple potential explanations for this: (1) the nonlinearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training (MacAvaney et al., 2022). We leave investigation of these avenues for future work.
@@ 6. Insights into Model Mechanisms
Towards a holistic approach in axiomatic analysis. Previous research investigating NRMs’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones (Câmara and Hauff, 2020). Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves. BERT’s “rediscovery” of BM25 indicates that it leverages a combination of retrieval heuristics to compute relevance, whereas previous analysis treats axioms independently. Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.
7. Conclusion
In this work, we mechanistically uncovered the core components of relevance scoring pathway of an NRM, revealing their ability to leverage contextual representations to perform a semantic-variant of BM25. Our granular analysis paves way for future directions. For example, future work could build on our IDF editing technique (§4.4) to adjust term importance, offering more control over model behavior and leading to improvements in model robustness against adversarial attacks, personalization, and bias mitigation. In addition, our evidence about the centrality of IDF information in the model’s embedding matrix raises the question of how crucial non-embedding weights are for task generalization, and future work could investigate whether fine-tuning only the embedding matrix with domain-specific IDF values can achieve robust performance while being far more parameter efficient. Finally, in this work, we analyze the behavior of a single cross-encoder in order to deeply analyze its mechanisms, and future work should investigate their generalizability to other neural IR models. Overall, identifying universal components or architecture-specific mechanisms will advance interpretability and inform the design of controllable and efficient neural ranking systems for real-world applications.

### v2

@@ 5.1 Formalizing the Hypothesized Function
  3. 3.
    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, modeling the product of IDF and TF in BM25.
By incorporating both the hypothesized computation function and the earlier components $U_{0}$ and $MS$, the linear model effectively represents the hypothesized circuit.
5.2 Assessing the Linear Model Fit
@@
We test whether our hypothesis holds by comparing the linear model’s effectiveness against the cross-encoder’s actual relevance scores.
@@
First, we train a linear regression model using our base dataset, limiting queries to five tokens to keep the number of coefficients manageable. For each forward pass, we extract two features for each query token: (1) the MS (Matching Score), calculated as the sum of the query token’s attention over document tokens from the 13 Matching Heads, and (2) its value along the top singular vector in $U_{0}$ of the decomposed embedding matrix. These features form the input x to the linear regression model, while the cross-encoder’s relevance scores are the target y. Finally, we evaluate how well the linear representation of the Semantic Scoring Circuit predicts the cross-encoder’s relevance scores with an 80/20 train-test split.
@@
The linear regression model achieves a high Pearson correlation (r = 0.8157, p < 0.001) with ground-truth relevance scores, showing that it captures the core of the cross-encoder’s scoring mechanism in a simplified and interpretable form. This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p < 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components effectively capture the signals that the cross-encoder utilizes for ranking. The linear model’s strong generalization to unseen datasets and varying query lengths (details in Appendix §E) further confirms that our circuit understanding captures the cross-encoder’s core mechanism for relevance computation.
6 Discussion
A two-stage process for relevance computation. Similar to Tenney et al. (2019), who find that BERT rediscovers the classical NLP pipeline, encoding syntax in its lower layers and semantics in its upper, our circuit analysis reveals a corresponding two-stage process. In the early and middle layers, Matching Heads extract lexical signals (e.g., soft-TF matches), reflecting prior work that argues document representations are query-dependent (Qiao et al., 2019). In the upper layers, Contextual Query Representation and Relevance Scoring Heads aggregate these signals into BM25-like relevance scores, consistent with Zhan et al. (2020), who argue that document tokens are largely query-independent. This dual-stage mechanism reconciles these two conflicting viewpoints by showing how query-specific signals migrate from document to query representations across layers and provides a more nuanced understanding of how relevance is computed.
Controllable IR. In §4.4, we show how model editing can be used to upscale or downscale term importance by modifying the encoded IDF values in the embeddings, effectively introducing “tuning dials” for fine-grained control over model behavior. This opens the door to personalization, where models could be adapted to user- or domain-specific preferences without full retraining.
Beyond personalization, these insights enable two additional downstream applications (preliminary experiments in Appendix §F). First, targeted model editing could mitigate unsafe or adversarial content without impairing the model’s ranking effectiveness. Second, aligning internal representations with ground-truth IDF scores could serve as a powerful initialization strategy for fine-tuning, leading to more efficient retrieval pipelines.
7 Conclusion
In this work, we mechanistically uncover the core components of the relevance scoring pathway in a BERT-based IR model, revealing how it leverages contextual representations to implement a semantic variant of BM25. Our fine-grained analysis lays the foundation for applying interventions to build more transparent and controllable models, enabling personalization, bias mitigation, and parameter-efficient adaptation. More broadly, identifying universal components or architecture-specific mechanisms contributes to the interpretability of IR models and informs the design of controllable, efficient ranking systems for real-world deployment.
Limitations
In this work, we focus on analyzing the behavior of a single cross-encoder in order to deeply analyze its mechanisms. Future work should investigate the generalizability of this behavior to other neural IR models with the same or different architectural base.
Additionally, while the Semantic Scoring circuit bears a strong resemblance to BM25, there may also be other factors influencing relevance computation that are out of the scope of investigation for this work. In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022). However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
Furthermore, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores. There are a couple potential explanations for this: (1) the non-linearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training MacAvaney et al. (2022). We leave the investigation of these avenues for future work.
Finally, although we intend for our analysis to inform future work on efficiency, transparency, and performance improvements, like most research, it is possible that a malicious actor could apply these insights to produce IR models that are robust to adversarial use cases.
Acknowledgment
We thank the Health NLP Lab for valuable suggestions and discussions that improved this work.
Appendix A Detailed Path Patching Methodology
Figure 7: Path patching methodology. Left: There are four forward passes: (1,2) Run model on baseline and perturbed inputs and cache activations. (3) To measure the effect of an upstream sender component (S) on a downstream receiver component (R), run the model on the baseline input, patch in S, freeze all other components, and cache the activation of R. (4) Run the model on the baseline input and patch in R. Right: Alternative visualization of step (3) using the residual stream. By allowing only the downstream receiver R to be recomputed when the sender S is patched, we effectively isolate the direct path from S to R, while preserving all other paths to R as they were in the baseline run.
The effect of a patch is measured by the difference in logits (which in the case of cross-encoders, is equivalent to the difference in relevance scores). This algorithm is then iteratively repeated for each important upstream component. For more information, we refer the reader to Wang et al. (2022) and Goldowsky-Dill et al. (2023).
Appendix B Activation Patching on Attention Heads
Prior to path patching, we conduct an exploratory analysis to compare the similarity between exact and semantic matches. Figure 8 shows that the patching effect has a high correlation of 0.96, suggesting that the model employs highly similar components for both exact and soft matches. These initial results on the indirect effect of these components provide a foundation to investigate their direct effects further with path patching.
Appendix C Detailed Path Patching Results
C.1 Path Patching to Final Logits and Intermediate Heads
We perform path patching on 2000 random query and document pairs from the STMC1 and TFC1 diagnostic datasets and collect the results.
In this section, we present the results of path patching on TFC1 and STMC1. First, we patch to the final logits and identify the shared Relevance Scoring Heads (Figure 9). Next, we patch to these heads and discover the shared Query Contextualization Heads (Figures 1011). Finally, by patching to the Query Contextualization Heads, we identify the shared Matching Heads (Figure 12).
Figure 8: Activation Patching Results for TFC and STMC diagnostic datasets shows a correlation of 0.96, which suggests that the model employs highly similar components for exact and soft matches.
Figure 9: Path Patching results from all attention heads to the final logits. From this, we identify that 10.1, 10.4, 10.7, 10.19, those causing > 30 % increase in ranking score, as most significant heads and focus our analysis on these four heads.
Figure 10: Path Patching results from upstream attention heads to Relevance Scoring Heads.
Figure 11: Path Patching results from upstream attention heads to Relevance Scoring Heads.
Figure 12: Path Patching results from upstream attention heads to Query Contextualization heads.
C.2 Path Patching to Matching Heads
We path patch the residual stream (resid_pre) on each of the Matching Heads using 100 random pairs from the TFC1 and STMC1 datasets (Figures 13, 14, 15, 16). Path patching to the Matching Heads, aimed at tracking the flow of soft-TF information, reveals that at the beginning of the residual stream at layer 0, which corresponds to the embeddings, has the most direct influence on the Matching Heads. In contrast, there is minimal impact from both the attention and MLP layers. This supports the conclusion that the Matching Heads are the primary “generator” of soft-TF, attending to soft matches based on semantic similarity from the embeddings. Thus, we conclude our tracking of soft-TF information, with the Matching Heads identified as the most upstream heads.
Figure 13: TFC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
Figure 14: TFC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
Figure 15: STMC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
Figure 16: STMC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
Appendix D Additional Details on Matching Heads and Additional Signals
In this section, we provide additional details on the analysis presented in §4.3.
Further Patching Experiment to Confirm Matching Heads Are Not Just Writing a Binary Signal of Existence. The signals written by most Matching Heads increase as the number of duplicate tokens rises, rather than remaining constant, proving that these signals are discrete rather than binary. To demonstrate this, we design an activation patching experiment (§4.3) that patches only the Matching Heads using the TFC2 Diagnostic Dataset. Figure 17 shows that for most Matching Heads, increasing the number of duplicate tokens results in a monotonically increasing pattern in their output signals. This behavior further supports our conclusion that Matching Heads compute and encode soft-TF signals rather than merely relaying binary information.
Figure 17: For the majority of Matching Heads, after the initial occurrence of $X_{\text{selected}}$, which causes a sharp increase in attention, subsequent occurrences result in only minimal incremental increases. This aligns with TFC2, which states that additional term occurrences yield smaller improvements.
Figure 18: Attention trends for selected terms are shown, with unselected terms following a similar pattern: longer documents generally increase the average attention assigned to all document tokens, regardless of whether they match a query token.
Matching Score Contains Term Saturation and Document Length Signals. The average soft-TF correlation score for Matching Heads (0.500) indicates that their attention values from query tokens to document tokens capture not only semantic similarity but also additional signals. To examine whether these attention values capture term saturation or document length effects, we analyze their behavior under using two diagnostic datasets: (1) TFC2 (increasing occurrences of a query term), and (2) LNC1 dataset (increasing number of irrelevant sentences to simulate the isolated effect of longer document length).
As the occurrences of $x_{\text{selected}}$ increase, its average attention value grows (Figure 17). On the other hand, the attention for $x_{\text{others}}$ remains relatively constant, which aligns with our soft-TF understanding: only document tokens that are semantically similar to a query token get nonzero attention values, proportional to the extent and frequency of semantic similarity.
@@ Appendix D Additional Details on Matching Heads and Additional Signals
Table 2: Comparison of Pearson, Spearman Rank, and NDCG@10 for SemanticBM and Baselines.
Furthermore, for most heads, the TFC2 trend is preserved in the attention values. For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 18). Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes. The varying degrees of influence on the heads’ attention values shown in Figure 17 and 18 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects. Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
Appendix E Generalization across IR datasets
We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept. Now, we ask the question: how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on? In this section, we extend our analysis to 12 different IR datasets [^2] Due to computational resources, we use a subset of BEIR datasets Thakur et al. (2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains. and investigate various query lengths.
Dataset. Since our linear model relies on fixed features based on query length, we stratify the dataset based on query length. To ensure sufficient samples, we aggregate samples from 12 datasets and only retain groups with more than 500 samples. We incorporate groups with query lengths from 1 to 22, which corresponds to the 99.95th percentile of MSMARCO’s query length distribution to exclude outliers with excessively long queries that the model has rarely seen during training. This process results in a total of 19 groups, comprising 278194 samples, with an average group size of 14641.790 samples. While this approach may be unconventional for retrieval, our goal is to simply demonstrate consistency between the linear model’s relevance score and the cross-encoder to confirm the hypothesized function of the Relevance Scoring Heads.
Experiments and Results. We train 22 linear regression models, one for each query length (1-22), to predict the cross-encoder’s relevance scores. We evaluate each model using: (1) Pearson Correlation: quantifies the correlation between predicted and actual relevance scores; (2) Spearman Rank Correlation: assesses the consistency of the ranked lists between predictions and cross-encoder outputs; (3) NDCG@10: measures ranking effectiveness and ensures no significant effectiveness discrepancies. We include BM25 and a randomized set of linear regression features as baselines.
Table 2 shows the results, and we report median values to account for observable skewness caused by outliers in the data. We observe a Pearson correlation of ranking scores with a median of 0.8401 (median p < 0.001), a Spearman rank correlation of 0.7619 (median p = 0.072), and an 88.4% alignment with cross-encoder effectiveness in terms of NDCG@10.
The experimental results confirm the hypothesized function of the Relevance Scoring Heads (§4.1). Since this linear model summarizes the whole circuit as it is structured to incorporate both the computation and the necessary components, the high correlation with the cross-encoder’s effectiveness shows that our circuit understanding has captured the core part of the cross-encoder’s relevance ranking mechanism.
Appendix F SVD Model Editing on Term Importance
In §4.4, we introduce a model editing technique to dynamically control the importance signal of query tokens, and here we present encouraging evidence to use the model editing technique for group-editing the importance signals in downstream tasks.
F.1 Mitigating Adversarial Attacks
Figure 19: Left: Decreasing the IDF of dangerous tokens shows an increasing trend in adversarial mitigation proportion. Right: The intervention maintains a high NDCG score at all levels, demonstrating that the general ranking capability of the cross-encoder remains unaffected.
First, we show the potential of the model editing in mitigating unsafe documents without disrupting the cross-encoder’s ranking performance, unlike retraining or regularization methods that often degrade it.
First, we construct a dataset using obscene and offensive words (LDNOOBW, 2015), filtering out multi-word entries and injecting these unsafe tokens into the safe samples of our TFC datasets. Next, we inspect the subgroup of queries and documents where inserting an unsafe token in the document significantly increases the relevance score. Our goal is to “erase” the effect of the dangerous token by reducing its importance. Among 17,537 adversarial samples, our approach achieves a 80.396% success rate (i.e., the unsafe document is ranked lower than the safe document). For example, when we downweight the target unsafe term by a large factor (e.g., -1200), the model maintains a high nDCG across all ranks, with a score of 0.9861, which is only a 1.39% drop in ranking performance. These results provide preliminary evidence for the potential of localized editing in the IDF-storing low-rank matrix.
F.2 Parameter Efficient Fine-Tuning
Our mechanistic analysis of the model’s relevance mechanisms reveals that term importance (IDF) is encoded in the embedding matrix. Since IDF determines how query terms are weighted in BM25, we hypothesize that only editing or fine-tuning the embedding matrix could provide comparable performance to fine-tuning the entire model, thus providing efficiency benefits.
To test this hypothesis with a preliminary experiment, we use the nfcorpus data, a Nutrition Fact retrieval dataset, from BEIR Thakur et al. (2021). We fine-tune on 10 epochs and evaluate performance across four conditions: (A1) full fine-tuning, (A2) IDF-aligned full fine-tuning (aligning token IDFs to the dataset using the model-editing method from §4.4), (B1) embedding-only fine-tuning, and (B2) IDF-aligned embedding-only fine-tuning. Results across three random (Figure 20 seeds show that IDF-aligned full fine-tuning (A2) achieves the highest NDCG@1, while IDF-aligned embedding-only fine-tuning (B2) outperforms unaligned embedding-only fine-tuning (B1). Additionally, full fine-tuning (A1) and embedding-only fine-tuning (B1) show no significant difference, supporting the idea that adaptation primarily refines the embedding matrix rather than the full model weights. These findings reveal preliminary indications that fine-tuning just on the embedding matrix could potentially be sufficient to adapt cross-encoders to new domains and reveal IDF-aligned initialization as a possible pathway for efficient adaptation.
Figure 20: For random seed 15, aligned full fine-tuning achieves the best performance, and aligned embeddings only achieve performance compared to full fine-tuning.
Appendix G Additional Discussion
@@
Towards a holistic approach in axiomatic analysis. Previous research investigating neural IR models’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones Câmara and Hauff (2020). Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves. BERT’s “rediscovery” of BM25 indicates that it leverages a combination of retrieval heuristics to compute relevance, whereas previous analysis treats axioms independently. Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.